OpenStudy (anonymous):

Im a little younger than the rest of you and i need help with wayy smaller situations but if there is someone out there who is wiling to help me please answer this question: How do you turn a mixed number to a decimal ?

OpenStudy (anonymous):

Quick question: do you know how to turn a fraction into a decimal (like \[1/3 = 0.333333333... or 3/4=0.75\] ). If yes, then turning a mixed number into a decimal is the same idea. Let's say the mixed number is \[3 ((4/5))\] (where 3 is the whole number and 4/5 is the fraction). Then we know we will already know what the number is to the left of the decimal point: 3 Now for the numbers to the left of the decimal point we look at the fraction: 4/5. If you know how to divide then you can set up the problem to \[4 \div 5\]. This will equal 0.8/. So if we combine 3+0.8=3.8. If you want to double check this you can also turn the mixed number into an improper fraction. So my example will look like 19/5 as an improper fraction. Again if you use division you will find that \[19 \div 5\] = 3.8. Does that answer your question?

OpenStudy (anonymous):

no not even close im in 5th grade

OpenStudy (anonymous):

im dealing with decimals mixed numbers and fractions

OpenStudy (anonymous):

Okay do you know division?

OpenStudy (anonymous):

If not, I would recommend getting a pencil and paper to write your problem out.

OpenStudy (anonymous):

I Know Division

OpenStudy (anonymous):

Wonderful. Just wondering, what is one of the problems you have that is a mixed number? That way we can work on it step by step.

OpenStudy (anonymous):

Umm... I Dont Really Have One Its Just A Subject we are working on in school You can make up A PROBLEM IF THATS OKAY WITH YOU sorry fot the caps lock

OpenStudy (anonymous):

for**

OpenStudy (anonymous):

Okay here is a mixed number: 4 5/6

OpenStudy (anonymous):

(where 4 is the whole number and 5/6 is the fraction).

OpenStudy (anonymous):

4and5 6th

OpenStudy (anonymous):

yess i know that part

OpenStudy (anonymous):

like a circle cut into 6 slices and only 5 shaded with 4 whole

OpenStudy (anonymous):

Excellent. So if we want to turn this into a decimal we need to figure out the numbers to the left of the decimal first. Numbers to the left of a decimal are whole numbers. What will be our whole number for this problem?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Okay so all that is left is to find the numbers that will be on the decimal side. So since you have already address that 4 is a whole number, we now want to focus on the fraction 5/6. Now you will need to do some division (if you want to use paper): so \[5 \div 6\]. Tell me what you get when you do long division.

OpenStudy (anonymous):

1 remainder 1

OpenStudy (anonymous):

0.83333

OpenStudy (anonymous):

Awesome. So now we just need to add our whole number with our fraction/decimal you just found. So 4+0.83333333= ????

OpenStudy (anonymous):

4.83333

OpenStudy (anonymous):

And that is the answer!

OpenStudy (anonymous):

I dont think i can go to school and put = 4.83333

OpenStudy (anonymous):

Most of the time a teacher will make you round your decimal. Besides, this is just the process you can take to solve the problem.

OpenStudy (anonymous):

so can i take some of the 3's off to make it 4.83 but than again i still dont completely Understand How we got this answer

OpenStudy (anonymous):

If your teacher says to round to two decimal places, then yes you can write it like 4.83. How about this; I'll give you a problem and you work on it (going through the process again), and tell me what you think the answer is. Then if you did something wrong I can see what part you may not understand.

OpenStudy (anonymous):

How about this: 6 3/8

OpenStudy (anonymous):

6.375 is the answer but thats from the calculator

OpenStudy (anonymous):

i dont get it myself

OpenStudy (anonymous):

Try not to use your calculator.

OpenStudy (anonymous):

i did the i got 6.2something

OpenStudy (anonymous):

So you can see where they got the 6 from in 6.375, right?

OpenStudy (anonymous):

I think you may be messing up when you are trying to do the division for the fraction.

OpenStudy (anonymous):

i know that the whole # always goes on the left side of the decimal i just need help when i divide on the right side what do i do with the remainder?

OpenStudy (anonymous):

Which number did you place inside the division sign: 3 or 8?

OpenStudy (anonymous):

3 d sign 8 *d sign means division sign*

OpenStudy (anonymous):

So you when you're doing long division you need to put 3 inside the division symbol and 8 on the outside. Since 8 can not go into 3, you need to add a decimal point and a 0 after the 3. So it will look like 3.0 divided by 8. So how many times does 8 go into 30?

OpenStudy (anonymous):

ooooo so 8 goes into 30 3 times and that will be 24

OpenStudy (anonymous):

So on top of the division sign you will have (so far) 0.3. So what is 30-24? And how many times does 8 go into that number?

OpenStudy (anonymous):

Ignore the second question actually

OpenStudy (anonymous):

30-24=6 and 8 can not go into 6 oh you said dont answer it

OpenStudy (anonymous):

Tis fine, but its good that you realize that 8 cannot go into 6. Anyway so after you subtract 30-24 you get 6; if you carry a zero down, you get 60. So how many times does 8 go into 60?

OpenStudy (anonymous):

8 goes in to 60 7 times that equals 56

OpenStudy (anonymous):

Mhm. So now the number on top of your dvision sign will be 0.37 (so far). So 60-56=???

OpenStudy (anonymous):

can U be a fan of me please i did it for u

OpenStudy (anonymous):

You should finish the problem first.

OpenStudy (anonymous):

60-56=4

OpenStudy (anonymous):

than will U be a fan?

OpenStudy (anonymous):

When you carry down a 0 you have 40. How many times does 8 go into 40?

OpenStudy (anonymous):

8 goes into 40 5 times and that equals exactly 40

OpenStudy (anonymous):

Then on top you have 0.375. And since 40-40=0 you are done with the division. So now you can add your decimal with the whole number: 6+.375=6.375. Does that make more sense?

OpenStudy (anonymous):

yes thank you so much no i really understand it and just a couple questions what country are you from is this a female or male and how old are you?

OpenStudy (anonymous):

I'm female, but I do not like to give out personal information. I would recommend practicing these problems (they become easier the more you do them).

OpenStudy (anonymous):

thats not much of personal information just what country and age there is millions of people in the same country and can be the same age

OpenStudy (anonymous):

All that matters though is that you got help for you problem correct? It doesn't really matter what country I am from or how old I am.

OpenStudy (anonymous):

it doesnt but i would like to know

OpenStudy (anonymous):

never mind its okay i dont want you to feel uncomfortable if im asking a question to personal im sorry dont feel the need to answer me or anyone else if you feel uncomfortable with any questions